OpenStudy (george.g101):

Which of the following is the simplified form ofthe seventh root of x times the seventh root of x times the seventh root of x ? A.) x to the 3 over seventh power B.) x to the 1 over seventh power C.) x to the 3 over 21st power D.) The 21st root of x

OpenStudy (george.g101):

Please please help @hartnn @agent0smith @YanaSidlinskiy

OpenStudy (george.g101):

I'm lost D=

OpenStudy (george.g101):

@jigglypuff314 @imqwerty

OpenStudy (kinged):

Okay, use this:\[x ^{\frac{ 1 }{ 7 }}\] is what seventh root of x means.

OpenStudy (kinged):

So, you just multiply them using the formula I previously supplied.

OpenStudy (george.g101):

Okk so x \[x\frac{ 1 }{ 7} * x\frac{ 1 }{ 7} * x\frac{ 1 }{ 7}* x\frac{ 1 }{ 7}* \]

OpenStudy (george.g101):

Wait so it would be C?

OpenStudy (george.g101):

No B?

OpenStudy (agent0smith):

Don't guess. Show work.

OpenStudy (kinged):

What is 1/7 + 1/7 +1/7?

OpenStudy (agent0smith):

And also make the question clear. How many x's are there? First there was three, now all of a sudden there's four.

OpenStudy (robtobey2):

\[\sqrt[7]{x}^3=x^{3/7} \]